### v1

# Regular Vaidya solutions of effective gravitational theories

Abstract
We initiate the study of the dynamics of spherically symmetric spacetimes beyond general relativity through exact solutions of the field equations of second-order effective gravitational theories defined solely in terms of the symmetries of the problem in 4 or more dimensions. We prove the existence of regular Vaidya solutions that represent the formation or change in mass of regular black holes (or black hole mimickers) due to the gravitational collapse or accretion of radiation, with the time-reverse of these processes being also included as solutions. Our treatment and results are remarkable in their simplicity and the possibilities they provide for novel theoretical applications and discoveries in the context of regular gravitational collapse. For instance, we illustrate how modifications of the gravitational dynamics can cure curvature singularities through what can be interpreted as an energy transfer between matter and gravitational degrees of freedom. As another application, we show how to incorporate spacetimes frequently used in the literature, designed geometrically to represent the formation and disappearance of regular black holes, as fully dynamical solutions within this formalism.
I Introduction
Black holes are one of the most fascinating theoretical constructs in gravitational physics. Each of their characteristic elements are examples of extreme behaviors of spacetime curvature due to gravitational effects: photon spheres illustrate the extreme bending of light into bounded orbits Luminet:1979nyg ; Perlick:2021aok , event horizons demonstrate the eternal capture of matter and energy inside black hole interiors Rindler:1956yx ; Gourgoulhon:2008pu , and singularities indicate the unbounded growth of gravitational fields and the consequent breakdown of general relativity itself Penrose:1964wq ; Senovilla:2014gza .
While observations have thus far shown an excellent agreement with the predictions of general relativity for the spacetime in the exterior of these gravitational objects LIGOScientific:2016aoc ; Cahillane:2022pqm ; EventHorizonTelescope:2019dse ; EventHorizonTelescope:2022wkp , numerous conceptual difficulties plague the description of their interior. Namely, the aforementioned existence of singularities Penrose:1964wq ; Senovilla:2014gza , the information loss problem Unruh:2017uaw ; Buoninfante:2021ijy , and the instability of inner horizons Simpson:1973ua ; Poisson1989 ; Poisson1990 ; Ori:1991zz .
It is widely expected that a full solution of these difficulties requires unifying quantum field theory and general relativity. While existing theories of quantum gravity are still far from providing a complete picture Carballo-Rubio:2025fnc , the leading low-energy effects of such theories can often be described in terms of effective geometric modifications of general relativity (recent reviews of gravitational collapse in quantum gravity include Eichhorn:2022bgu ; Ashtekar:2023cod ; Platania:2023srt ). Research along this line has, however, stumbled upon substantial difficulties. There is evidence that regularizing general black hole interiors requires field equations of order higher than second Knorr:2022kqp ; Borissova:2023kzq , but higher-order field equations pose important challenges even for basic working notions such as well posedness Kovacs:2020pns ; Kovacs:2020ywu ; Brady:2023dgu .
This work presents a breakthrough in the study of the effective theories describing black holes. We show for the first time the existence of simple exact solutions representing the formation of regular black holes for a wide class of second-order gravitational theories in 4 or more dimensions. The starting point of the analysis is the definition of the space of second-order theories compatible with the symmetries of the problem, with no further restrictions. We then select the matter source to search for Vaidya solutions in this general space, and determine the subset of these theories for which these solutions are singularity-free.
In general relativity, the Vaidya solution is a generalization of the Schwarzschild solution with null dust (radiation) as a matter source which modifies the mass of the central black hole Vaidya:1951fdr ; Vaidya:1953zza ; Vaidya:1966zza , or its time-reversal in which the central object is instead a white hole. Due to its simplicity and physical significance, this key solution is used extensively in discussions of black hole physics across different levels, from textbooks and lecture notes on the subject Strominger:1994tn ; Poisson:2009pwt to research papers discussing core aspects of black hole physics, including diverse definitions of black hole boundaries Schnetter:2005ea ; Ben-Dov:2006zmw ; Bengtsson:2008jr ; Nielsen:2010wq and the phenomena of mass inflation Ori:1991zz ; Frolov:2017rjz ; Bonanno:2020fgp ; Carballo-Rubio:2021bpr ; Carballo-Rubio:2024dca and black hole evaporation Parikh:1998ux ; Hayward:2005gi ; Frolov:2014jva . Showing that an equally simple generalization of this famous solution exists for second-order effective gravitational theories, including cases describing dynamical regular black holes instead, is a crucial development in the study of physics beyond general relativity.
II Second-order effective gravitational actions
In the following, we focus our attention on spherically symmetric situations. While our main focus is working in $D=4$ dimensions, the formalism introduced below can describe in a unified manner the dynamics of spacetimes for any dimensionality $D\geq 4$. Keeping the dimensionality general will also be useful in the discussion of some of the examples below. We can thus consider a general metric of the form
$g^{(D)}_{\mu\nu}(y)\text{d}y^{\mu}\text{d}y^{\nu}=g_{ab}(x)\text{d}x^{a}\text{d}x^{b}+R^{2}(x)\text{d}\Omega_{(D-2)}^{2},$ (1)
where $D$–dimensional coordinates are labeled by $\{y^{\mu}\}_{\mu=0}^{D-1}$, $R(x)$ is a scalar field on the 2–dimensional space with coordinates $\{x^{a}=x^{0},x^{1}\}$ (and will itself be used as one of these coordinates later on), and $\text{d}\Omega_{(D-2)}^{2}$ is the line element on the unit $(D-2)$–sphere.
The dynamics of the $D$–dimensional metric $g^{(D)}_{\mu\nu}$ and any other matter fields is described by a $D$–dimensional action $\mathscr{S}^{(D)}=\int\text{d}^{D}y\sqrt{-g^{(D)}}\left[\mathscr{L}_{\rm G}^{(D)}/\kappa+\mathscr{L}_{\rm M}^{(D)}\right]$, where $\mathscr{L}_{\rm G}^{(D)}$ and $\mathscr{L}_{\rm M}^{(D)}$ are the gravitational and matter Lagrangian densities, respectively, and $\kappa=16\pi$ in natural units. The equations of motion can be obtained by a variational principle and then restricted to spherically symmetric metrics of the form in Eq. (1). However, due to the compact nature of the rotation group, it is equivalent to restrict the $D$–dimensional actions to spherically symmetric metrics of the form in Eq. (1) and then perform variations Palais:1979rca ; Fels:2001rv ; Frausto:2024egp . This is the route we will follow here.
When restricting the action $\mathscr{S}^{(D)}$ to spherically symmetric metrics, the resulting gravitational Lagrangian is $\Omega_{(D-2)}\mathscr{L}_{\rm G}$, where $\Omega_{(D-2)}=\int\text{d}\Omega_{(D-2)}=2\pi^{(D-1)/2}/\Gamma(D/2-1/2)$ is a numerical constant and $\mathscr{L}_{\rm G}$ a function of a 2–dimensional metric $g_{ab}(x)$ and a scalar field $R(x)$, corresponding to general 2–dimensional dilaton (2DD) gravities Colleaux:2019ckh . In general, the equations of motion will be of arbitrarily high order, depending on the specific $D$–dimensional theory considered. This space of theories is completely general and includes the dynamics of the spherically symmetric sector of any $D$–dimensional metric theory that admits a variational principle formulation.
Within the general space of 2DD gravities, there exists a subset with second-order field equations. This is a corollary of the seminal work of Horndeski (see, e.g., Horndeski:1974wa ; Kobayashi:2019hrl ), particularized to a 2–dimensional metric $g_{ab}$, which implies that the most general Lagrangian density satisfying this requirement can be written as:
$\mathscr{L}_{\rm G}=H_{2}(R,X)-H_{3}(R,X)\square R+H_{4}(R,X)\mathcal{R}-2\partial_{X}H_{4}(R,X)\left[(\square R)^{2}-\nabla^{a}\nabla^{b}R\nabla_{a}\nabla_{b}R\right],$ (2)
where $\mathcal{R}$ is the Ricci scalar of $g_{ab}$, $\{H_{i}(R,X)\}_{i=1}^{4}$ are generic functions of two variables, and we have defined $X=\left(\nabla R\right)^{2}$. Let us note that, in higher dimensions, the most general Horndeski Lagrangian contains terms proportional to an additional function $H_{5}(R,X)$, but these terms vanish identically for 2–dimensional metrics.
These 2–dimensional Horndeski (2DH) gravities can be understood either as describing the dynamics of gravity coupled to a matter scalar field in 2–dimensional spacetimes, or the dynamics of gravity in $D$–dimensional spherically symmetric spacetimes (or spacetimes with other symmetries which allow for an equivalent reduction). This second interpretation is the one relevant for us. Instead of committing to specific families of gravitational theories, we are exploiting the symmetries of the problem to make calculations tractable, which allows us to prove general results and distinguishes our approach from recent works Bueno:2024eig ; Bueno:2024zsx ; Bueno:2025gjg ; Fernandes:2025fnz .
The spherically symmetric reduction of general relativity without a cosmological constant corresponds to the particular case
$H_{2}=(D-2)(D-3)R^{D-4},\quad H_{3}=(D-2)R^{D-3},\quad H_{4}=R^{D-2}.$ (3)
In the following we will only consider theories that asymptote to general relativity at large distances, recovering these expressions in the $R\rightarrow\infty$ (weak gravity) limit.
Expressions for the variations of the Lagrangian density in Eq. (2) are known Kunstatter:2015vxa ; Colleaux:2019ckh . Adding matter, we can write down the following three independent field equations:
$\beta\nabla_{a}\nabla_{b}R-g_{ab}\left(\frac{1}{2}\alpha+\beta\square R\right)+\left(\partial_{X}\alpha-\partial_{R}\beta\right)\nabla_{a}R\nabla_{b}R=8\pi T_{ab},$ (4)
where we have defined
$\alpha=H_{2}+X\partial_{R}\left(H_{3}-2\partial_{R}H_{4}\right),\quad\beta=X\partial_{X}\left(H_{3}-2\partial_{R}H_{4}\right)-\partial_{R}H_{4},$ (5)
as well as the conserved $2$–dimensional source
$T_{ab}=R^{D-2}T^{(D)}_{\mu\nu}\delta^{\mu}_{a}\delta^{\nu}_{b},$ (6)
with the $D$–dimensional stress-energy tensor $T^{(D)}_{\mu\nu}$ defined using the standard Hilbert prescription.
Solutions of the field equations (4) in vacuum ($T_{ab}=0$) are known (e.g., Ziprick:2010vb ; Kunstatter:2015vxa ). However, to the best of our knowledge, it is the first time that exact solutions in the presence of a source, describing dynamical black holes, are studied in this formalism.
We note for completeness that, due to its generality, this formalism can provide a natural framework to discuss theories with a varying gravitational coupling, an implementation of which has been recently discussed in the context of the gravitational collapse of timelike dust Bonanno:2023rzk . Establishing this connection requires a dedicated analysis that will be performed elsewhere.
III Vaidya solutions
In the rest of the paper we will be considering $D$–dimensional metrics of the form
$\text{d}s^{2}=-f(V,R)\text{d}V^{2}+2\text{d}V\text{d}R+R^{2}\text{d}\Omega^{2}_{D-2},$ (7)
where $f(V,R)$ is a function to be determined by solving the field equations, and $V$ is an ingoing null coordinate. Some geometric relations that will be useful later are $\nabla_{V}\nabla_{V}R=-\left(f\partial_{R}f-\partial_{V}f\right)/2$, $\nabla_{V}\nabla_{R}R=\nabla_{R}\nabla_{V}R=\partial_{R}f/2$, $\nabla_{R}\nabla_{R}R=0$, $(\nabla R)^{2}=f$ and $\square R=\partial_{R}f$. All calculations below apply, with just a few sign changes, to the time-reversal of the metric (7), by implementing the transformation $V\rightarrow-V\propto U$, where $U$ would be an outgoing null coordinate.
The $D$–dimensional Vaidya solution Iyer:1989nd is a solution of the Einstein field equations with matter source
$T^{(D)}_{\mu\nu}=\frac{(D-2)\dot{M}(V)}{8\pi R^{D-2}}\partial_{\mu}V\partial_{\nu}V,$ (8)
where $\dot{M}(V)=\text{d}M(V)/\text{d}V$. We will now describe how the existence of solutions with this same matter source is a feature of more general 2DH gravities beyond general relativity.
The three independent components of Eq. (4) yield the following equations:
$\partial_{X}\alpha-\partial_{R}\beta$ $=0,$ (9)
$\alpha+\beta\partial_{R}f$ $=0,$ (10)
$\beta\partial_{V}f$ $=2(D-2)\dot{M}(V).$ (11)
The solution to these equations is given by a function $\Theta\left(R,X\right)$ satisfying the off-shell relations
$\partial_{R}\Theta(R,X)=\alpha(R,X),\quad\partial_{X}\Theta(R,X)=\beta(R,X),$ (12)
as well as the on-shell relation
$\left.\Theta\left(R,X\right)\right|_{X=f(V,R)}=2(D-2)M(V).$ (13)
Specific solutions can be generated in two ways, either by specifying functions $\alpha(R,X)$ and $\beta(R,X)$ satisfying the integrability condition (9) and finding $\Theta(R,X)$ and $f(V,R)$, or by specifying a function $f(V,R)$ and deriving $\Theta(R,X)$, $\alpha(R,X)$ and $\beta(R,X)$.
All Vaidya solutions discussed here are of type D in the Petrov–Pirani–Penrose classification Petrov:2000bs ; Pirani:1956wr ; Newman:1961qr ; Penrose:1985bww , as is the case in general relativity Stephani:2003tm . The principal double null directions $\{N^{(i)}\}_{i=1}^{2}$, satisfying $C_{\mu\nu\rho[\sigma}N^{(i)}_{\gamma]}N^{(i)\nu}N^{(i)\rho}=0$ where $C_{\mu\nu\rho\sigma}$ is the Weyl tensor, are the same as for general relativity (see, e.g., Coudray:2021pjr ), namely $N^{(1)}=\partial/\partial R$ and $N^{(2)}=\partial/\partial V+\left(f/2\right)\partial/\partial R$.
The analysis above includes any metric gravitational theory with second-order field equations in spherical symmetry, recovering known results for theories such as Lovelock gravity Kobayashi:2005ch ; Maeda:2005ci ; Dominguez:2005rt ; Ghosh:2008jca ; Cai:2008mh but also proving the existence of previously unknown Vaidya solutions in other frameworks such as quasitopological gravities Myers:2010ru ; Oliva:2010eb ; Bueno:2016xff ; Hennigar:2017ego ; Ahmed:2017jod ; Bueno:2019ycr ; Bueno:2022res ; Moreno:2023rfl , which will be discussed in the examples below. It can also accommodate Vaidya geometries obtained in studies of effective quantum gravity corrections, regardless of whether these are singular or regular Borissova:2022mgd ; Delaporte:2024and . While in the following we focus on regular geometries, this wide range of potential applications is interesting in its own right.
IV Regular Vaidya solutions
The degree of regularity of any geometry of the form in Eq. (1) can be checked by inspecting the components of curvature tensors in locally regular coordinate systems Rinne:2005df (see also  Carballo-Rubio:2019fnb ). For the metric in Eq. (7), regularity in terms of finiteness of curvature scalars at the origin amounts to two conditions on the redshift function, $f|_{R=0}=1$ and $\partial_{R}f|_{R=0}=0$. These conditions imply, together with Eqs. (10) and  (11), the following on-shell divergent behaviour in the $R\rightarrow 0$ limit:
$\alpha(R,f)\propto R^{p},\quad\beta(R,f)\propto R^{p-1},\quad\text{with}\quad p\leq-1.$ (14)
The on-shell function $\Theta(R,f)$ does not depend on $R$, as shown in Eq. (13), and it is equivalent to the mass function defined in Kunstatter:2015vxa .
Due to dimensional reasons, satisfying these conditions requires the introduction of at least one new length scale $\ell$, when compared to the singular counterparts of these geometries in general relativity. Depending on the relative value of $\ell$ (which we use as a placeholder for the potentially multiple new parameters in the solution) and $M(V)$ at a given moment of time, the corresponding geometry describes either a regular black hole or a horizonless object. Typically, for a given value of $\ell$, there is a maximum threshold of mass, determined kinematically, above which the corresponding geometries describe regular black holes. This threshold can be calculated explicitly for specific examples, such as the Hayward case discussed below (see, e.g., Carballo-Rubio:2022nuj for details).
Note that the value of the new parameter $\ell$ is not fixed a priori in our approach. Choosing macroscopic values of $\ell$ can thus provide a phenomenological description of the formation of black hole mimickers (e.g. gravastar-like objects Mazur:2001fv ; Mazur:2004fk ; Visser:2003ge ) from the collapse of null dust, as well as of radiation emission and absorption processes which change the mass of such objects. The study of black hole mimickers is an important area of research in its own right that has gained traction recently (see recent reviews in Carballo-Rubio:2025fnc ; Bambi:2025wjx ). Providing models for the dynamical formation and evolution of such objects is one of the key open challenges in this field, which our results can be directly applied to.
V Examples
We start noting that, for general relativity, $\alpha=(D-2)(D-3)R^{D-4}(1-X)$ and $\beta=-(D-2)R^{D-3}$. These functions satisfy Eq. (9) and the corresponding Vaidya metric is generated by $\Theta(R,X)=(D-2)R^{D-3}(1-X)$, which results in $f(V,R)=1-2M(V)/R^{D-3}$.
The first example are Ziprick-Kunstatter gravities, defined as the $D$–dimensional generalization of the 4-dimensional theories studied by Ziprick and Kunstatter Ziprick:2010vb (see also Kunstatter:2015vxa ; Barenboim:2024dko ; Barenboim:2025ckx ), with $H_{2}=\text{d}^{2}\Phi(R)/\text{d}R^{2}$, $H_{3}=\text{d}\Phi(R)/\text{d}R$ and $H_{4}=\Phi(R)$, with a a function $\Phi(R)$ behaving as $\Phi(R)\simeq R^{D-2}$ for $R\rightarrow\infty$ [thus recovering general relativity at large distances]. We thus have $\alpha=\text{d}^{2}\Phi/\text{d}R^{2}(1-X)$, $\beta=-\text{d}\Phi/dR$ which, in accordance with the regularity conditions discussed above, implies that $\Phi(R)$ must diverge at least as $\Phi(R)\propto R^{-1}$ for $R\rightarrow 0$. We also have $\Theta(R,X)=(1-X)\text{d}\Phi/\text{d}R$ and $f(V,R)=1-2(D-2)M(V)/\left(\text{d}\Phi/\text{d}R\right)$.
The second example are quasitopological gravities, higher-curvature theories defined for $D\geq 5$ and displaying second-order field equations on spherically symmetric backgrounds Myers:2010ru ; Oliva:2010eb ; Bueno:2016xff ; Hennigar:2017ego ; Ahmed:2017jod ; Bueno:2019ycr ; Bueno:2022res ; Moreno:2023rfl , which are thus included in our analysis. To show this explicitly, we note that the two cases above are particular instances of a general expression $\Theta(R,X)=(D-2)R^{D-1}\Psi(Z)$, where $\Psi(Z)$ is a function of the variable $Z=(1-X)/R^{2}$. We then obtain $\alpha=(D-2)R^{D-2}\left[(D-1)\Psi-2Z\text{d}\Psi/\text{d}Z\right]$ and $\beta=-(D-2)R^{D-3}\text{d}\Psi/\text{d}Z$. These expressions coincide with the ones obtained when considering the spherically symmetric reduction of quasitopological gravities Bueno:2024eig ; Bueno:2024zsx ; Bueno:2025gjg .
The third example will be constructed following the inverse approach, starting from the $D$–dimensional generalization of the Hayward metric Hayward:2005gi , $f(V,R)=1-2R^{2}M(V)/\left[R^{D-1}+2\ell^{D-2}M(V)\right]$. Using this expression along with Eq. (13), we can determine that $\Theta(R,X)=(D-2)(1-X)R^{D-1}/\left[R^{2}-(1-X)\ell^{D-2}\right]$, from which $\alpha$ and $\beta$ can be obtained. This case can also be written in terms of a function $\Psi(Z)=Z/(1-\ell^{D-2}Z)$. The same approach can be applied to any regular black hole metric.
VI Singularity regularization and energy transfer
The matter source in Eq. (8) is singular at $R=0$. However, the modified Vaidya solutions we consider here are explicitly regular by construction, whenever the conditions in Eq. (14) are satisfied.
A conceptual understanding of this issue is provided by the following splitting of the gravitational field equations:
$G^{(D)}_{\mu\nu}$ $=8\pi\left[T^{(D)}_{\mu\nu}+\mathcal{T}^{(D)}_{\mu\nu}\right],$ (15)
where $G^{(D)}_{\mu\nu}$ is the $D$–dimensional Einstein tensor, $T^{(D)}_{\mu\nu}$ is the matter source in Eq. (8), and $\mathcal{T}^{(D)}_{\mu\nu}$ is an effective stress-energy tensor coming from the terms regularizing the solutions to the Einstein equations. All these tensors are independently conserved.
For the regular Vaidya solutions, we have the following key expressions:
$G^{(D)}_{VV}$ $=-f(V,R)G^{(D)}_{VR}-\frac{(D-2)^{2}}{R}\frac{\dot{M}(V)}{\beta(R,f)},$
$G^{(D)}_{VR}$ $=\frac{(D-2)(D-3)}{2R^{2}}\left[f(V,R)-1\right]+\frac{(D-2)}{2R}\partial_{R}f(V,R),$
$8\pi\mathcal{T}^{(D)}_{VV}$ $=-f(V,R)G^{(D)}_{VR}-\frac{(D-2)^{2}}{R}\frac{\dot{M}(V)}{\beta(R,f)}-\frac{(D-2)\dot{M}(V)}{R^{D-2}}.$ (16)
These equations can be straightforwardly obtained by writing these components of the $D$–dimensional Einstein tensor in terms of variations of the gravitational action with respect to $g_{ab}$, and using explicitly Eqs. (3), (4) and (11). We can see that all the components of the Einstein tensor are regular. On the other hand, $\mathcal{T}^{(D)}_{VV}$ is singular and actually cancels the singularity in $T^{(D)}_{VV}$.
The only component of the matter stress-energy tensor $T^{(D)}_{\mu\nu}$ is an ingoing radial null flux $T^{(D)}_{VV}$ that is divergent at $R=0$ and has a constant integral over all spheres throughout its collapse, $\Omega_{(D-2)}T_{VV}=\Omega_{(D-2)}R^{D-2}T_{VV}^{(D)}$. On the other hand, the $\mathcal{T}^{(D)}_{VV}$ component of the effective stress-energy tensor contains an ingoing flux proportional to $\dot{M}$ that is also divergent at $R=0$, but with an overall integral over spheres that depends on the radius. Indeed, the piece proportional to $\dot{M}$ in $\Omega_{(D-2)}\mathcal{T}_{VV}=\Omega_{(D-2)}R^{D-2}\mathcal{T}_{VV}^{(D)}$ vanishes in the limit $R\rightarrow\infty$ and is switched on progressively as the radius decreases, until becoming $-\Omega_{(D-2)}T_{VV}$ exactly at $R=0$. Due to $\mathcal{T}^{(D)}_{\mu\nu}$ being conserved, this ingoing flux of energy must be created at the expense of leaving behind a residual distribution of energy, encoded in the remaining non-zero pieces of $\mathcal{T}^{(D)}_{\mu\nu}$, which describe how the energy in the collapsing or accreting null dust is stored in the gravitational field of the regular black hole after being effectively transferred.
The signs of these fluxes depend on the sign of $\dot{M}$. For $\dot{M}>0$ the flux in $T^{(D)}_{\mu\nu}$ has positive energy and the flux in $\mathcal{T}^{(D)}_{\mu\nu}$ has negative energy, while for $\dot{M}<0$ these signs are reversed. In other words, $\mathcal{T}^{(D)}_{\mu\nu}$ always acts as a regulator of $T^{(D)}_{\mu\nu}$ regardless of the sign of $\dot{M}(V)$. Depending on the evolution of $M(V)$, the corresponding geometries can describe processes of formation, accretion or disappearance of regular black holes. Combining both signs of $\dot{M}$ allows us to construct geometries with trapped regions but no event horizons (see Fig. 1).
Such geometric constructions have frequently been used to model the expected effects of semiclassical backreaction, and are ubiquitous in the discussions of the evaporation of regular black holes, including Hayward:2005gi ; Frolov:2014jva ; Frolov:2017rjz ; Cardoso:2023guh ; Carballo-Rubio:2024dca . Here these arise as exact solutions, which opens the possibility of a thorough understanding of the dynamical properties of evaporating regular black holes, such as their behavior under radial perturbations and the backreaction of the mass inflation instability Frolov:2017rjz ; Bonanno:2020fgp ; Carballo-Rubio:2021bpr ; Carballo-Rubio:2024dca . Adding semiclassical sources, along the lines of the recent analysis for Reissner–Nordström black holes Boyanov:2025otp (see also Barcelo:2020mjw ; Barcelo:2022gii ), would also allow for a self-consistent modeling of the semiclassical dynamics of these regular objects.
Figure 1: Left: Vaidya solution in general relativity representing the formation of a Schwarzschild black hole from the collapse of null dust originating from past null infinity, as encoded in the flux component of $T_{\mu\nu}^{(D)}$, represented with blue arrows and contained within the blue dashed lines. The trapped region, characterized by $f(V,R)<0$, is shaded in gray. Right: regular Vaidya solution representing the formation and disappearance of a regular black hole due to the collapse and accretion of null dust. Aside from the classical matter source $T^{(D)}_{\mu\nu}$, the effective source $\mathcal{T}^{(D)}_{\mu\nu}$ has two distinct components: an ingoing flux (red arrows) with energy of opposite sign to that of the null dust, and the energy distribution deposited in the regular black hole (red shaded region). Both classical matter and effective sources cancel identically at $R=0$. The energy stored in the regular black hole grows while $\dot{M}>0$ (in between the blue dashed lines), decreases while $\dot{M}<0$ (in between the green dashed lines), and remains constant otherwise.
VII Conclusions
In this paper, we have shown how the study of the field equations of general second-order effective gravitational theories describing spherically symmetric spacetimes can provide important insights into the geometric structure of black holes beyond general relativity. In particular, we have generalized the famous Vaidya solution of the Einstein-Hilbert theory to a wider class of gravitational actions, including theories in which the central curvature singularity is not formed. This result is remarkable due to its mathematical simplicity as well as its potential implications for the study of the effective low-energy dynamics of black holes in quantum gravity.
The singularity-free generalizations of the Vaidya solution discussed here maintain the simplicity that is characteristic of the original solution, while presenting new intriguing features that offer a novel window into the interplay of quantum effects and gravity. In particular, the energy in the collapsing null dust is transferred to gravitational degrees of freedom and stored in the resulting regular black hole. We hypothesize that this energy transfer may be a key aspect in the formation of macroscopic black holes in quantum gravity.
Another remarkable feature of our results is their generality. The starting point of our analysis of spherically-symmetric geometries encompasses any new (quantum) gravitational effects describable with second-order field equations in $D\geq 4$ dimensions. To our knowledge, it is the first time that progress in the study of the dynamics of regular black holes has been made in such a general framework. While quantum gravity may ultimately involve higher-derivative field equations, the results discussed here are a strong indication that second-order field equations may suffice to describe key dynamical features of regular black holes, and are thus worth exploring in detail.
In general relativity, exact solutions describing the formation of black holes paved the way for the development of numerical relativity, and the subsequent blossoming of the field of gravitational wave astronomy. It is our hope that our present results will provide a similar push toward the development of numerics beyond general relativity, this being a promising emergent area of research. Our methods and results provide a simple and powerful framework to go beyond the test-field approximation in the study of the gravitational effects of matter and information inside black holes, which may provide new insights into the fundamentals of gravitational physics.
Acknowledgements.
The authors are grateful to Jose Beltrán Jiménez for insightful discussions on 2–dimensional Horndeski gravities, and to Carlos Barceló, Vitor Cardoso and Stefano Liberati for feedback on the first version of the manuscript. RCR acknowledges financial support provided by the Spanish Government through the Ramón y Cajal program (contract RYC2023-045894-I), the Grant No. PID2023-149018NB-C43 funded by MCIN/AEI/10.13039/501100011033, and the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033, as well as the hospitality of the Center of Gravity, a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. VB acknowledges support from the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587, as well as from the Spanish Government through the Grants No. PID2020-118159GB-C43, PID2020-118159GB-C44, PID2023-149018NB-C43 and PID2023-149018NB-C44 (funded by MCIN/AEI/10.13039/501100011033).

### v2

@@
 Research along this line has, however, stumbled upon substantial difficulties.
-There is evidence that regularizing general black hole interiors requires field equations of order higher than second Knorr:2022kqp ; Borissova:2023kzq , but higher-order field equations pose important challenges even for basic working notions such as well posedness Kovacs:2020pns ; Kovacs:2020ywu ; Brady:2023dgu .
+There is evidence that regularizing general black hole interiors requires field equations of order higher than second Knorr:2022kqp ; Borissova:2023kzq , but higher-order field equations pose important challenges even for basic working notions such as well-posedness Kovacs:2020pns ; Kovacs:2020ywu ; Brady:2023dgu .
 This work presents a breakthrough in the study of the effective theories describing black holes.
@@
 In general relativity, the Vaidya solution is a generalization of the Schwarzschild solution with null dust (radiation) as a matter source which modifies the mass of the central black hole Vaidya:1951fdr ; Vaidya:1953zza ; Vaidya:1966zza , or its time-reversal in which the central object is instead a white hole.
-Due to its simplicity and physical significance, this key solution is used extensively in discussions of black hole physics across different levels, from textbooks and lecture notes on the subject Strominger:1994tn ; Poisson:2009pwt to research papers discussing core aspects of black hole physics, including diverse definitions of black hole boundaries Schnetter:2005ea ; Ben-Dov:2006zmw ; Bengtsson:2008jr ; Nielsen:2010wq and the phenomena of mass inflation Ori:1991zz ; Frolov:2017rjz ; Bonanno:2020fgp ; Carballo-Rubio:2021bpr ; Carballo-Rubio:2024dca and black hole evaporation Parikh:1998ux ; Hayward:2005gi ; Frolov:2014jva .
+Due to its simplicity and physical significance, this key solution is used extensively in discussions of black hole physics across different levels, from textbooks and lecture notes on the subject Strominger:1994tn ; Poisson:2009pwt to research papers discussing core aspects of black hole physics, including diverse definitions of black hole boundaries Schnetter:2005ea ; Ben-Dov:2006zmw ; Bengtsson:2008jr ; Nielsen:2010wq , the phenomenon of mass inflation Ori:1991zz ; Frolov:2017rjz ; Bonanno:2020fgp ; Carballo-Rubio:2021bpr ; Carballo-Rubio:2024dca , and black hole evaporation Parikh:1998ux ; Hayward:2005gi ; Frolov:2014jva .
 Showing that an equally simple generalization of this famous solution exists for second-order effective gravitational theories, including cases describing dynamical regular black holes instead, is a crucial development in the study of physics beyond general relativity.
@@
 where $D$–dimensional coordinates are labeled by $\{y^{\mu}\}_{\mu=0}^{D-1}$, $R(x)$ is a scalar field on the 2–dimensional space with coordinates $\{x^{a}=x^{0},x^{1}\}$ (and will itself be used as one of these coordinates later on), and $\text{d}\Omega_{(D-2)}^{2}$ is the line element on the unit $(D-2)$–sphere.
-The dynamics of the $D$–dimensional metric $g^{(D)}_{\mu\nu}$ and any other matter fields is described by a $D$–dimensional action $\mathscr{S}^{(D)}=\int\text{d}^{D}y\sqrt{-g^{(D)}}\left[\mathscr{L}_{\rm G}^{(D)}/\kappa+\mathscr{L}_{\rm M}^{(D)}\right]$, where $\mathscr{L}_{\rm G}^{(D)}$ and $\mathscr{L}_{\rm M}^{(D)}$ are the gravitational and matter Lagrangian densities, respectively, and $\kappa=16\pi$ in natural units.
+We start considering as motivation the dynamics of the $D$–dimensional metric $g^{(D)}_{\mu\nu}$ and any other matter fields, described by a $D$–dimensional action $\mathscr{S}^{(D)}=\int\text{d}^{D}y\sqrt{-g^{(D)}}\left[\mathscr{L}_{\rm G}^{(D)}/\kappa+\mathscr{L}_{\rm M}^{(D)}\right]$, where $\mathscr{L}_{\rm G}^{(D)}$ and $\mathscr{L}_{\rm M}^{(D)}$ are the gravitational and matter Lagrangian densities, respectively, and $\kappa=16\pi$ in natural units.
 The equations of motion can be obtained by a variational principle and then restricted to spherically symmetric metrics of the form in Eq.
@@
 (1) and then perform variations Palais:1979rca ; Fels:2001rv ; Frausto:2024egp .
-This is the route we will follow here.
-When restricting the action $\mathscr{S}^{(D)}$ to spherically symmetric metrics, the resulting gravitational Lagrangian is $\Omega_{(D-2)}\mathscr{L}_{\rm G}$, where $\Omega_{(D-2)}=\int\text{d}\Omega_{(D-2)}=2\pi^{(D-1)/2}/\Gamma(D/2-1/2)$ is a numerical constant and $\mathscr{L}_{\rm G}$ a function of a 2–dimensional metric $g_{ab}(x)$ and a scalar field $R(x)$, corresponding to general 2–dimensional dilaton (2DD) gravities Colleaux:2019ckh .
+When restricting the action $\mathscr{S}^{(D)}$ to spherically symmetric metrics, the resulting gravitational Lagrangian is $\Omega_{(D-2)}\mathscr{L}_{\rm G}$, where $\Omega_{(D-2)}=\int\text{d}\Omega_{(D-2)}=2\pi^{(D-1)/2}/\Gamma(D/2-1/2)$ is a numerical constant and $\mathscr{L}_{\rm G}$ a function of a 2–dimensional metric $g_{ab}(x)$ and a scalar field $R(x)$, corresponding to general 2–dimensional dilaton (2DD) gravities (where using the term dilaton stresses its geometric origin).
 In general, the equations of motion will be of arbitrarily high order, depending on the specific $D$–dimensional theory considered.
@@
 In the following we will only consider theories that asymptote to general relativity at large distances, recovering these expressions in the $R\rightarrow\infty$ (weak gravity) limit.
-Expressions for the variations of the Lagrangian density in Eq. (2) are known Kunstatter:2015vxa ; Colleaux:2019ckh . Adding matter, we can write down the following three independent field equations:
+Expressions for the variations of the Lagrangian density in Eq.
+(2) are known (see the detailed discussion in Colleaux:2019ckh ).
+For our purposes here, it suffices to write explicitly the following three independent field equations in the presence of matter:
 $\beta\nabla_{a}\nabla_{b}R-g_{ab}\left(\frac{1}{2}\alpha+\beta\square R\right)+\left(\partial_{X}\alpha-\partial_{R}\beta\right)\nabla_{a}R\nabla_{b}R=8\pi T_{ab},$ (4)
@@
 $\alpha=H_{2}+X\partial_{R}\left(H_{3}-2\partial_{R}H_{4}\right),\quad\beta=X\partial_{X}\left(H_{3}-2\partial_{R}H_{4}\right)-\partial_{R}H_{4},$ (5)
-as well as the conserved $2$–dimensional source
+as well as the $2$–dimensional source
 $T_{ab}=R^{D-2}T^{(D)}_{\mu\nu}\delta^{\mu}_{a}\delta^{\nu}_{b},$ (6)
-with the $D$–dimensional stress-energy tensor $T^{(D)}_{\mu\nu}$ defined using the standard Hilbert prescription.
-Solutions of the field equations (4) in vacuum ($T_{ab}=0$) are known (e.g., Ziprick:2010vb ; Kunstatter:2015vxa ).
-However, to the best of our knowledge, it is the first time that exact solutions in the presence of a source, describing dynamical black holes, are studied in this formalism.
+where the $D$–dimensional stress-energy tensor $T^{(D)}_{\mu\nu}$ is defined using the standard Hilbert prescription.
 We note for completeness that, due to its generality, this formalism can provide a natural framework to discuss theories with a varying gravitational coupling, an implementation of which has been recently discussed in the context of the gravitational collapse of timelike dust Bonanno:2023rzk .
@@
 Specific solutions can be generated in two ways, either by specifying functions $\alpha(R,X)$ and $\beta(R,X)$ satisfying the integrability condition (9) and finding $\Theta(R,X)$ and $f(V,R)$, or by specifying a function $f(V,R)$ and deriving $\Theta(R,X)$, $\alpha(R,X)$ and $\beta(R,X)$.
+When considering the particular cases in which $M(V)$ becomes a constant, we recover the vacuum solutions in Kunstatter:2015vxa .
+The most general Lagrangian density introduced in Kunstatter:2015vxa differs from Eq.
+(2) by a total derivative (see Colleaux:2019ckh ), which in particular implies that the theories defined by Eq.
+(2) must also satisfy Birkhoff’s theorem.
 All Vaidya solutions discussed here are of type D in the Petrov–Pirani–Penrose classification Petrov:2000bs ; Pirani:1956wr ; Newman:1961qr ; Penrose:1985bww , as is the case in general relativity Stephani:2003tm .
 The principal double null directions $\{N^{(i)}\}_{i=1}^{2}$, satisfying $C_{\mu\nu\rho[\sigma}N^{(i)}_{\gamma]}N^{(i)\nu}N^{(i)\rho}=0$ where $C_{\mu\nu\rho\sigma}$ is the Weyl tensor, are the same as for general relativity (see, e.g., Coudray:2021pjr ), namely $N^{(1)}=\partial/\partial R$ and $N^{(2)}=\partial/\partial V+\left(f/2\right)\partial/\partial R$.
-The analysis above includes any metric gravitational theory with second-order field equations in spherical symmetry, recovering known results for theories such as Lovelock gravity Kobayashi:2005ch ; Maeda:2005ci ; Dominguez:2005rt ; Ghosh:2008jca ; Cai:2008mh but also proving the existence of previously unknown Vaidya solutions in other frameworks such as quasitopological gravities Myers:2010ru ; Oliva:2010eb ; Bueno:2016xff ; Hennigar:2017ego ; Ahmed:2017jod ; Bueno:2019ycr ; Bueno:2022res ; Moreno:2023rfl , which will be discussed in the examples below.
+The analysis above includes any metric gravitational theory with second-order field equations in spherical symmetry, recovering known results for theories such as Lovelock gravity Kobayashi:2005ch ; Maeda:2005ci ; Dominguez:2005rt ; Ghosh:2008jca ; Cai:2008mh but also proving the existence of previously unknown Vaidya solutions in other frameworks such as quasitopological gravities Myers:2010ru ; Oliva:2010eb ; Bueno:2016xff ; Hennigar:2017ego ; Ahmed:2017jod ; Bueno:2019ycr ; Bueno:2022res ; Moreno:2023rfl ; Bueno:2024dgm , which will be discussed in the examples below.
 It can also accommodate Vaidya geometries obtained in studies of effective quantum gravity corrections, regardless of whether these are singular or regular Borissova:2022mgd ; Delaporte:2024and .
@@
 $\alpha(R,f)\propto R^{p},\quad\beta(R,f)\propto R^{p-1},\quad\text{with}\quad p\leq-1.$ (14)
-The on-shell function $\Theta(R,f)$ does not depend on $R$, as shown in Eq. (13), and it is equivalent to the mass function defined in Kunstatter:2015vxa .
+The on-shell function $\Theta(R,f)$ does not depend on $R$, as shown in Eq. (13).
 Due to dimensional reasons, satisfying these conditions requires the introduction of at least one new length scale $\ell$, when compared to the singular counterparts of these geometries in general relativity.
@@
 (9) and the corresponding Vaidya metric is generated by $\Theta(R,X)=(D-2)R^{D-3}(1-X)$, which results in $f(V,R)=1-2M(V)/R^{D-3}$.
-The first example are Ziprick-Kunstatter gravities, defined as the $D$–dimensional generalization of the 4-dimensional theories studied by Ziprick and Kunstatter Ziprick:2010vb (see also Kunstatter:2015vxa ; Barenboim:2024dko ; Barenboim:2025ckx ), with $H_{2}=\text{d}^{2}\Phi(R)/\text{d}R^{2}$, $H_{3}=\text{d}\Phi(R)/\text{d}R$ and $H_{4}=\Phi(R)$, with a a function $\Phi(R)$ behaving as $\Phi(R)\simeq R^{D-2}$ for $R\rightarrow\infty$ [thus recovering general relativity at large distances].
+The first example are the $D$–dimensional generalization of the theories defined for 4-dimensional spherically symmetric spacetimes by Ziprick and Kunstatter Ziprick:2010vb (see also Kunstatter:2015vxa ; Barenboim:2024dko ; Barenboim:2025ckx ), namely $H_{2}=\text{d}^{2}\Phi(R)/\text{d}R^{2}$, $H_{3}=\text{d}\Phi(R)/\text{d}R$ and $H_{4}=\Phi(R)$, with a a function $\Phi(R)$ behaving as $\Phi(R)\simeq R^{D-2}$ for $R\rightarrow\infty$ [thus recovering general relativity at large distances].
 We thus have $\alpha=\text{d}^{2}\Phi/\text{d}R^{2}(1-X)$, $\beta=-\text{d}\Phi/dR$ which, in accordance with the regularity conditions discussed above, implies that $\Phi(R)$ must diverge at least as $\Phi(R)\propto R^{-1}$ for $R\rightarrow 0$.
 We also have $\Theta(R,X)=(1-X)\text{d}\Phi/\text{d}R$ and $f(V,R)=1-2(D-2)M(V)/\left(\text{d}\Phi/\text{d}R\right)$.
-The second example are quasitopological gravities, higher-curvature theories defined for $D\geq 5$ and displaying second-order field equations on spherically symmetric backgrounds Myers:2010ru ; Oliva:2010eb ; Bueno:2016xff ; Hennigar:2017ego ; Ahmed:2017jod ; Bueno:2019ycr ; Bueno:2022res ; Moreno:2023rfl , which are thus included in our analysis.
-To show this explicitly, we note that the two cases above are particular instances of a general expression $\Theta(R,X)=(D-2)R^{D-1}\Psi(Z)$, where $\Psi(Z)$ is a function of the variable $Z=(1-X)/R^{2}$.
-We then obtain $\alpha=(D-2)R^{D-2}\left[(D-1)\Psi-2Z\text{d}\Psi/\text{d}Z\right]$ and $\beta=-(D-2)R^{D-3}\text{d}\Psi/\text{d}Z$.
-These expressions coincide with the ones obtained when considering the spherically symmetric reduction of quasitopological gravities Bueno:2024eig ; Bueno:2024zsx ; Bueno:2025gjg .
+The second example are quasitopological gravities Myers:2010ru ; Oliva:2010eb ; Bueno:2016xff ; Hennigar:2017ego ; Ahmed:2017jod ; Bueno:2019ycr ; Bueno:2022res ; Moreno:2023rfl ; Bueno:2024dgm , higher-curvature theories defined for $D\geq 5$ and displaying second-order field equations for spherically symmetric spacetimes, which are thus included in our analysis.
+To show this explicitly, we note that the two cases above are particular instances of a general expression $\Theta(R,X)=\Psi(Z)R^{2}\text{d}\Phi/\text{d}R$, where $\Psi(Z)$ is a function of the variable $Z=(1-X)/R^{2}$.
+For $\Phi(R)=R^{D-2}$, we obtain $\alpha=(D-2)R^{D-2}\left[(D-1)\Psi-2Z\text{d}\Psi/\text{d}Z\right]$ and $\beta=-(D-2)R^{D-3}\text{d}\Psi/\text{d}Z$.
+These expressions coincide with the ones obtained when considering the spherically symmetric reduction of quasitopological gravities for the study of the gravitational collapse of timelike thin shells Bueno:2024eig ; Bueno:2024zsx and timelike dust Bueno:2025gjg in $D\geq 5$ (embedding specific regular black holes in quasitopological gravities involves additional constraints that prevent specific metrics to be realized in some dimensionalities).
 The third example will be constructed following the inverse approach, starting from the $D$–dimensional generalization of the Hayward metric Hayward:2005gi , $f(V,R)=1-2R^{2}M(V)/\left[R^{D-1}+2\ell^{D-2}M(V)\right]$.
@@
 VI Singularity regularization and energy transfer
-The matter source in Eq. (8) is singular at $R=0$. However, the modified Vaidya solutions we consider here are explicitly regular by construction, whenever the conditions in Eq. (14) are satisfied.
+The matter source in Eq. (8) is singular at $R=0$. However, the corresponding Vaidya solutions are explicitly regular whenever the conditions in Eq. (14) are satisfied.
 A conceptual understanding of this issue is provided by the following splitting of the gravitational field equations:
@@
 On the other hand, $\mathcal{T}^{(D)}_{VV}$ is singular and actually cancels the singularity in $T^{(D)}_{VV}$.
-The only component of the matter stress-energy tensor $T^{(D)}_{\mu\nu}$ is an ingoing radial null flux $T^{(D)}_{VV}$ that is divergent at $R=0$ and has a constant integral over all spheres throughout its collapse, $\Omega_{(D-2)}T_{VV}=\Omega_{(D-2)}R^{D-2}T_{VV}^{(D)}$.
+The only component of the matter stress-energy tensor $T^{(D)}_{\mu\nu}$ is an ingoing radial null flux $T^{(D)}_{VV}$ that is divergent at $R=0$, and its integral $\Omega_{(D-2)}T_{VV}=\Omega_{(D-2)}R^{D-2}T_{VV}^{(D)}$ is constant over all spheres throughout its collapse.
 On the other hand, the $\mathcal{T}^{(D)}_{VV}$ component of the effective stress-energy tensor contains an ingoing flux proportional to $\dot{M}$ that is also divergent at $R=0$, but with an overall integral over spheres that depends on the radius.
@@
 Such geometric constructions have frequently been used to model the expected effects of semiclassical backreaction, and are ubiquitous in the discussions of the evaporation of regular black holes, including Hayward:2005gi ; Frolov:2014jva ; Frolov:2017rjz ; Cardoso:2023guh ; Carballo-Rubio:2024dca .
-Here these arise as exact solutions, which opens the possibility of a thorough understanding of the dynamical properties of evaporating regular black holes, such as their behavior under radial perturbations and the backreaction of the mass inflation instability Frolov:2017rjz ; Bonanno:2020fgp ; Carballo-Rubio:2021bpr ; Carballo-Rubio:2024dca .
+Here these can be constructed as exact solutions, which opens the possibility of a thorough understanding of the dynamical properties of evaporating regular black holes, such as their behavior under radial perturbations and the backreaction of the mass inflation instability Frolov:2017rjz ; Bonanno:2020fgp ; Carballo-Rubio:2021bpr ; Carballo-Rubio:2024dca .
 Adding semiclassical sources, along the lines of the recent analysis for Reissner–Nordström black holes Boyanov:2025otp (see also Barcelo:2020mjw ; Barcelo:2022gii ), would also allow for a self-consistent modeling of the semiclassical dynamics of these regular objects.
-Figure 1: Left: Vaidya solution in general relativity representing the formation of a Schwarzschild black hole from the collapse of null dust originating from past null infinity, as encoded in the flux component of $T_{\mu\nu}^{(D)}$, represented with blue arrows and contained within the blue dashed lines.
+Recent related numerical breakthroughs can be found in Barenboim:2024dko ; Barenboim:2025ckx .
+Figure 1: Left: Vaidya solution in general relativity representing the formation of a Schwarzschild black hole from the collapse of null dust originating from past null infinity, as encoded in the flux component of $T_{\mu\nu}^{(D)}$, indicated by blue arrows and contained within the blue dashed lines.
 The trapped region, characterized by $f(V,R)<0$, is shaded in gray.
@@
 Acknowledgements.
-The authors are grateful to Jose Beltrán Jiménez for insightful discussions on 2–dimensional Horndeski gravities, and to Carlos Barceló, Vitor Cardoso and Stefano Liberati for feedback on the first version of the manuscript.
+The authors are grateful to Jose Beltrán Jiménez for insightful discussions on 2–dimensional Horndeski gravities, Carlos Barceló, Vitor Cardoso and Stefano Liberati for feedback on the first version of the manuscript, and Ángel Murcia for useful correspondence.
 RCR acknowledges financial support provided by the Spanish Government through the Ramón y Cajal program (contract RYC2023-045894-I), the Grant No.